# Information Limits of Joint Community Detection and Finite Group Synchronization

### v1

@@ 1 Introduction
Finding this limit enables us to benchmark the performance of existing algorithms and assess their potential for improvement. Notably, the optimal algorithm that maximizes the probability of exact recovery points to the maximum a posteriori (MAP) estimation. Meanwhile, since there is no preference regarding the underlying community assignment and group elements, the MAP estimation is equivalent to the maximum likelihood estimation (MLE) under a uniform prior. Therefore, the performance of MLE serves as the information-theoretic limits such that if MLE fails in exact recovery with high probability as the network size $n$ grows, no algorithm, regardless of its complexity, can succeed with high probability.
Our contribution
@@
    which agrees with the existing result of information limit on the SBM [1, 3, 50, 54].
@@
    By examining the performance of existing efficient algorithms [23, 6, 12, 66, 24], our theory demonstrates a significant performance gap between the information limit and those approaches (see Figure 2), suggesting a considerable room of improvement exists.
@@
Organization
The rest of this paper is organized as follows: In the remaining of Section 1, we introduce some additional related works and the notations for our analysis. Then, in Section 2 we formulate the MLE program for recovery on our statistical model and define the evaluation measurement. Next, in Section 3 we present our theory on the information-theoretic limits, and Section 4 is devoted to discussions on the result. We end up with the proofs in Section 5, and leave the technical details to the Appendix.
@@ 1.1 Related works
On the side of group synchronization, the goal is to recover the underlying group elements from a set of noisy pairwise measurements. A large body of literature focuses on the aforementioned random rewiring model (e.g. [62, 26, 21, 48, 65, 14, 13, 40]) which independently replaces each pairwise clean measurement with some random element. The fundamental information limits for exact recovery are generally studied in [14]. Regarding the algorithms, similar to the development of community detection, various approaches are developed for recovery depending on the specific group of interest, including semidefinite relaxation [61, 40], spectral methods [61, 5, 11, 56, 59, 34], and message passing [60], along with many theoretical investigations on their performances [61, 70, 48, 40, 49, 33, 68].
The joint problem of community detection and group synchronization is an emerging research topic (e.g. [26, 21, 6, 45]) motivated by recent scientific applications such as the cryo-electron microscopy single-particle reconstruction [29, 62, 69, 22] aforementioned. In particular, the statistical model adopted in this paper was initially proposed in [24], which also studied exact recovery by SDP with the performance guarantee provided. Recently, several spectral method-based approaches [23, 12, 66] are designed for efficient recovery. As there is a lack of indicators on the performance of these algorithms, the information-theoretic limit obtained in this work fills in the blanks.
@@ 1.2 Notations
Here, we define a series of notations for our statistical model that will be frequently used. Given the network of size $n$ with $K$ underlying communities, we denote $\bm{\kappa}^{*}:=\{\kappa_{1}^{*},\ldots,\kappa_{n}^{*}\}$ as the set of true community assignment, and $\bm{g}^{*}:=\{g_{1}^{*},\ldots,g_{n}^{*}\}$ as the set of true group elements. Similarly, the estimated community assignments and group elements are denoted as $\bm{\kappa}:=\{\kappa_{1},\ldots,\kappa_{n}\}$ and $\bm{g}:=\{g_{1},\ldots,g_{n}\}$, respectively. We use $n_{k}^{*}$ and $S_{k}^{*}$ to denote the size of the $k$-th community and the set of nodes belonging to it, respectively, i.e. $S_{k}^{*}:=\{i|\kappa_{i}=k\}$ and $n_{k}^{*}:=|S_{k}^{*}|$ for $k=1,\ldots K$, and we define $S_{k}:=\{i|\kappa_{i}=k\}$ and $n_{k}:=|S_{k}|$ analogous to $S_{k}^{*}$ and $n_{k}^{*}$. Then, we denote $\bm{x}^{*}:=\{\bm{\kappa}^{*},\bm{g}^{*}\}$ and $\bm{x}:=\{\bm{\kappa},\bm{g}\}$ as the true and estimated parameters. By denoting $\mathcal{E}$ as the set of edge connections and $\mathfrak{G}:=\{g_{ij}|(i,j)\in\mathcal{E}\}$ as the set of group transformations observed on each edge, we define $\bm{y}:=\{\mathcal{E},\mathfrak{G}\}$ as the whole observation of the network. Lastly, given any set of nodes $S$, we define $\mathcal{E}(S)=\{(i,j)|i,j\in S,(i,j)\in\mathcal{E}\}$ as the set of edge connections with $S$. For any two sets of nodes $S_{1}$ and $S_{2}$, we denote $\mathcal{E}(S_{1},S_{2})=\{(i,j)|i\in S_{1},j\in S_{2},(i,j)\in\mathcal{E})\}$ as the set of edges connecting $S_{1}$ and $S_{2}$. We summarize these notations in Table 1.
@@ 2.1 Maximum likelihood estimation
The derivation of (1) is deferred to Section 5. Here, we assume the cluster sizes $\{n_{k}^{*}\}_{k=1}^{K}$ are given and therefore the constraint $n_{k}=n_{k}^{*}$ ensures our recovery is valid.
@@ 2.2 Evaluation measurement
where $\Pi_{K}$ denotes the permutation group on $\{1,\ldots,K\}$. Notably, $\Pi_{K}$ is introduced to eliminate the ambiguity on the permutation of cluster labels. That is, any permutation on the labels does not affect the underlying partition, and thus the evaluation metric should be invariant to permutations.
The same idea applies to the metric for group elements: one can only identify them within each community up to some global offset $\bar{g}\in\mathcal{G}_{M}$, as there is no basis to distinguish $g_{i}$ with $g_{i}\bar{g}$ given the pairwise transformations. Therefore, similar to (5), we define
@@
In other words, $\mathcal{P}_{e,\text{c}}(\psi)$ (resp. $\mathcal{P}_{e,\text{g}}(\psi)$) stands for the maximum probability of giving a wrong estimation $\bm{\kappa}$ (resp. $\bm{g}$) by $\psi$ over all possible ground truth $\bm{x}^{*}$. We remark that in our model setting, the error probability does not depend on the specific choice of $\bm{x}^{*}$ as there is no preference. Therefore, one should expect both $\mathcal{P}\left(\textrm{dist}_{\text{c}}\left(\bm{\kappa},\bm{\kappa}^{*}\right)\neq 0\mid\bm{x}^{*}\right)$ and $\mathcal{P}\left(\textrm{dist}_{\text{g}}\left(\bm{g},\bm{g}^{*}\right)\neq 0\mid\bm{x}^{*}\right)$ are invariant to $\bm{x}^{*}$.
@@ 3 Main results
This section presents the information-theoretic limits for exact recovery on the statistical model defined in Section 1. Notably, even though we aim to recover both the (i) communities and (ii) group elements simultaneously, there might exist scenarios where the recovery of one is possible but impossible for the another. Because of this, we study the conditions for recovering (i) and (ii) separately. Specifically, for each of them, we provide
@@
  • Information-theoretic lower bound, below which the MLE (1) fails to recover with a probability bounded away from zero
In the rest of our analysis, we focus on the scenario of having two communities of equal sizes, and the order of group $M$ is assumed to be a fixed integer as $n$ grows.
@@ 3.1 Exact recovery of communities
$\frac{a+b}{2}-(1+o(1))\sqrt{\frac{ab}{M}}>1+c$ (8)
then the MLE (1) exactly recovers the communities w.p. at least $1-n^{-c}$, i.e. $\mathcal{P}_{e,\mathrm{c}}(\psi_{\textup{MLE}})\leq n^{-c}$.
@@
$\frac{a+b}{2}-\sqrt{\frac{ab}{M}}<1$ (9)
then the MLE (1) fails to recover the communities w.p. at least $\frac{3}{5}$, i.e. $\mathcal{P}_{e,\mathrm{c}}(\psi_{\textup{MLE}})\geq\frac{3}{5}$.
Proof strategy
The proofs of Theorems 3.1 and 3.2 are deferred to Section 5.2 and Section 5.3, respectively. Notably, our proof technique for the upper bound significantly differs from those traditional approaches employed in community detection or group synchronization (e.g. [1, Theorem 2] for SBM and [14, Theorem 1] for synchronization) in the way of handling different hypotheses. To be concrete, since the MLE fails when there exists a wrong hypothesis $\bm{x}=\{\bm{\kappa},\bm{g}\}$ with a higher likelihood than the ground truth $\bm{x}^{*}=\{\bm{\kappa}^{*},\bm{g}^{*}\}$, i.e.
In addition, our proof relies on several sharp analyses related to random graph theory (see e.g. Theorem A.4 in Appendix A), which serves as one of the main technical challenges. We leave some discussion in Section 4.
@@
$\frac{a+b}{2}-\sqrt{\frac{ab}{M}}=1$ (12)
@@ 3.2 Exact recovery of group elements
then the MLE (1) exactly recovers the group elements w.p. at least $1-n^{-c+o(1)}$, i.e. $\mathcal{P}_{e,\mathrm{g}}(\psi_{\textup{MLE}})\leq n^{-c+o(1)}$.
@@
then the MLE (1) fails to recover the group elements w.p. at least $0.45$, i.e. $\mathcal{P}_{e,\mathrm{g}}(\psi_{\textup{MLE}})\geq 0.45$.
The proofs of Theorems 3.3 and 3.4 are deferred to Section 5.4 and Section 5.5 respectively, which are simply extended from the ones of Theorems 3.1 and 3.2. Similar to the threshold (12), Theorems 3.3 and 3.4 suggest a phase transition exists for the exact recovery of group elements, given as

### v2

@@ 1 Introduction
Finding this limit enables us to benchmark the performance of existing algorithms and assess their potential for improvement. Notably, the joint problem for recovery can be viewed as a channel decoding problem from an information theory perspective (similar to the argument made in [2, Page 4]). Specifically, the input to the channel consists of the $n$ nodes that encodes the clustering labels and group elements, while the channel outputs correspond to the (${n\choose 2}$ pairwise observations – each representing whether a connection exists between a node pair and, if so, the associated group transformation observation. In this way, the algorithm that maximizes the probability of correctly recovering the communities and the group elements is the Maximum A Posteriori (MAP) decoder. Moreover, since the community assignments are assumed to be uniformly distributed, the MAP decoder coincides with the Maximum Likelihood (ML) decoder. Therefore, the performance of MLE serves as the information-theoretic limits such that if MLE fails in exact recovery with high probability as the network size $n$ goes to infinity, no algorithm, regardless of its complexity, can succeed with high probability.
1.1 Our contribution
@@
    which agrees with the existing result of information-theoretic limit on the SBM [2, 4, 53, 57].
@@
    By examining the performance of existing efficient algorithms [24, 7, 13, 70, 25], our theory demonstrates a significant performance gap between the information-theoretic limit and those approaches (see Figure 2), suggesting a considerable room of improvement exists (see Section 4 for a detailed discussion).
@@
1.2 Organization
The rest of this paper is organized as follows: In the remaining of Section 1, we introduce some additional related works and the notations for our analysis. Then, in Section 2 we formulate the MLE program for recovery on our statistical model and define the evaluation measurement. Next, in Section 3 we present our theory on the information-theoretic limits, and Section 4 is devoted to discussions on the result. We end up with proofs of the theorems in the remaining sections, from Section 6 for Theorem 3.1, to Section 9 for Theorem 3.4. We leave the technical details to the Appendix.
@@ 1.3 Related works
On the side of group synchronization, the goal is to recover the underlying group elements from a set of noisy pairwise measurements. A large body of literature focuses on the aforementioned random rewiring model (e.g. [66, 27, 22, 51, 69, 15, 14, 41]) which independently replaces each pairwise clean measurement with some random element. The fundamental information-theoretic limits for recovery are generally studied in several prior works including [15, 61, 71] regardless of the specific choice of group. Moreover, [43] focuses on SO(2) (angular synchronization) and U(1) group and provides asymptotic thresholds of recovery for several popular estimators including minimum MSE (Bayes Optimal Estimator) and MLE. Regarding the algorithms, similar to the development of community detection, various approaches are developed for recovery depending on the specific group of interest, including semidefinite relaxation [65, 41], spectral methods [65, 6, 12, 59, 63, 35], and (approximate) message passing [64, 61] (which is conjectured to achieve the information-theoretic limits), along with many theoretical investigations on their performances [65, 75, 51, 41, 52, 34, 73].
The joint problem of community detection and group synchronization is an emerging research topic (e.g. [27, 22, 7, 47]) motivated by recent scientific applications such as the cryo-electron microscopy single-particle reconstruction [30, 66, 74, 23] aforementioned. In particular, the statistical model adopted in this paper was initially proposed in [25], which also studied exact recovery by SDP with the performance guarantee provided. Recently, several spectral method-based approaches [24, 13, 70] are designed for efficient recovery. As there is a lack of indicators on the performance of these algorithms, the information-theoretic limit obtained in this work fills in the blanks. A detailed discussion on the existing algorithms and the comparison between their performances with the information-theoretic limit, is provided in Section 4.
@@ 1.4 Notations
Here, we define a series of notations for our statistical model that will be frequently used. Given the network of size $n$ with $K$ underlying communities, we denote $\bm{\kappa}^{*}:=\{\kappa_{1}^{*},\ldots,\kappa_{n}^{*}\}$ as the set of true community assignment, and $\bm{g}^{*}:=\{g_{1}^{*},\ldots,g_{n}^{*}\}$ as the set of true group elements. Similarly, the estimated community assignments and group elements are denoted as $\bm{\kappa}:=\{\kappa_{1},\ldots,\kappa_{n}\}$ and $\bm{g}:=\{g_{1},\ldots,g_{n}\}$, respectively. We use $n_{k}^{*}$ and $S_{k}^{*}$ to denote the size of the $k$-th community and the set of nodes belonging to it, respectively, i.e.,$S_{k}^{*}:=\{i|\kappa_{i}=k\}$ and $n_{k}^{*}:=|S_{k}^{*}|$ for $k=1,\ldots K$, and we define $S_{k}:=\{i|\kappa_{i}=k\}$ and $n_{k}:=|S_{k}|$ analogous to $S_{k}^{*}$ and $n_{k}^{*}$. Then, we denote $\bm{x}^{*}:=\{\bm{\kappa}^{*},\bm{g}^{*}\}$ and $\bm{x}:=\{\bm{\kappa},\bm{g}\}$ as the true and estimated parameters. By denoting $\mathcal{E}$ as the set of edge connections and $\mathfrak{G}:=\{g_{ij}|(i,j)\in\mathcal{E}\}$ as the set of group transformations observed on each edge, we define $\bm{y}:=\{\mathcal{E},\mathfrak{G}\}$ as the whole observation of the network. Lastly, given any set of nodes $S$, we define $\mathcal{E}(S)=\{(i,j)|i,j\in S,(i,j)\in\mathcal{E}\}$ as the set of edge connections with $S$. For any two sets of nodes $S_{1}$ and $S_{2}$, we denote $\mathcal{E}(S_{1},S_{2})=\{(i,j)|i\in S_{1},j\in S_{2},(i,j)\in\mathcal{E})\}$ as the set of edges connecting $S_{1}$ and $S_{2}$. We summarize these notations in Table 1.
@@ 2.1 Maximum likelihood estimation
The derivation of (1) is deferred to Section 5.2. Here, we assume the cluster sizes $\{n_{k}^{*}\}_{k=1}^{K}$ are given and therefore the constraint $n_{k}=n_{k}^{*}$ ensures our recovery is valid.
@@ 2.2 Evaluation measurement
where $\Pi_{K}$ denotes the permutation group on $\{1,\ldots,K\}$. Notably, $\Pi_{K}$ is introduced to eliminate the ambiguity on the permutation of cluster labels. That is, any permutation on the labels does not affect the underlying partition, and thus the evaluation metric should be invariant to permutations. The same idea applies to the metric for group elements: one can only identify them within each community up to some global offset $\bar{g}\in\mathcal{G}_{M}$, as there is no basis to distinguish $g_{i}$ with $g_{i}\bar{g}$ given the pairwise transformations. Therefore, similar to (5), we define
@@
In other words, $\mathcal{P}_{e,\text{c}}(\psi)$ (resp. $\mathcal{P}_{e,\text{g}}(\psi)$) stands for the maximum probability of giving a wrong estimation $\bm{\kappa}$ (resp. $\bm{g}$) by $\psi$ over all possible ground truth $\bm{x}^{*}$.
We remark that in our model setting, the error probability does not depend on the specific choice of $\bm{x}^{*}$ as there is no preference. Therefore, one should expect both $\mathcal{P}\left(\textrm{dist}_{\text{c}}\left(\bm{\kappa},\bm{\kappa}^{*}\right)\neq 0\mid\bm{x}^{*}\right)$ and $\mathcal{P}\left(\textrm{dist}_{\text{g}}\left(\bm{g},\bm{g}^{*}\right)\neq 0\mid\bm{x}^{*}\right)$ are invariant to $\bm{x}^{*}$
@@ 3 Main results
This section presents the information-theoretic limits for exact recovery on the statistical model defined in Section 1. We focus on the scenario of having two communities ($K=2$) of equal sizes ($n_{1}=n_{2}$), and the order of group $M$ is assumed to be a fixed integer as $n$ grows. Notably, even though we aim to recover both the (i) communities and (ii) group elements simultaneously, there might exist scenarios where the recovery of one is possible but impossible for the another. Because of this, we study the conditions for recovering (i) and (ii) separately. Specifically, for each of them, we provide
@@
  • Information-theoretic lower bound, below which the MLE (1) fails to exactly recover with a probability bounded away from zero
@@ 3.1 Exact recovery of communities
$\frac{a+b}{2}-(1+o(1))\sqrt{\frac{ab}{M}}>1+c,$ (8)
then the MLE (1) exactly recovers the communities w.p. at least $1-n^{-c}$, i.e.,$\mathcal{P}_{e,\mathrm{c}}(\psi_{\textup{MLE}})\leq n^{-c}$.
@@
$\frac{a+b}{2}-\sqrt{\frac{ab}{M}}<1,$ (9)
then the MLE (1) fails to exactly recover the communities w.p. at least $\frac{3}{5}$, i.e., $\mathcal{P}_{e,\mathrm{c}}(\psi_{\textup{MLE}})\geq\frac{3}{5}$.
@@
$\frac{a+b}{2}-\sqrt{\frac{ab}{M}}=1,$ (10)
@@
Proof strategy
The proofs of Theorems 3.1 and 3.2 are deferred to Section 6 and Section 7, respectively. Notably, our proof technique for the upper bound significantly differs from those traditional approaches employed in community detection or group synchronization (e.g. [2, Theorem 2] for SBM and [15, Theorem 1] for synchronization) in the way of handling different hypotheses. To be concrete, since the MLE fails when there exists a wrong hypothesis $\bm{x}=\{\bm{\kappa},\bm{g}\}$ with a higher likelihood than the ground truth $\bm{x}^{*}=\{\bm{\kappa}^{*},\bm{g}^{*}\}$, i.e.,
In addition, our proof relies on several sharp analyses related to random graph theory (see e.g., Theorem A.4 in Appendix A), which serves as one of the main technical challenges. We leave some discussion in Section 4.3.
An intuitive justification for Theorem 3.1
Here, we provide an intuitive justification for the upper bound (Theorem 3.1). Notably, the analysis here is intentionally informal as it involves a few heuristic assumptions and approximations. To begin with, consider the regime where $Mp(1-q)>q(1-p)$. In this case, the MLE (1) maximizes inner-cluster edges while maintaining group transformation consistency. For a specific node $i$, let us define $N^{(i)}_{\text{inner}}\sim\text{Binom}\left(\frac{n}{2}-1,p\right)$ and $N^{(i)}_{\text{inter}}\sim\text{Binom}\left(\frac{n}{2},q\right)$ as the number of inner and inter-cluster connections with the node $i$, respectively. Then, the MLE (1) fails if there exists a node $i$, such that
$\underbrace{N^{(i)}_{\text{inter}}\cdot\mathbbm{1}\left(g_{ij}=g_{i}g^{-1}_{j},\;\forall j\text{ connected to }i,\text{ and }\kappa_{j}\neq\kappa_{i}\right)}_{=:\tilde{N}^{(i)}_{\text{inter}}}\geq N^{(i)}_{\text{inner}},$
for some group elements $\{g_{j}\}_{j=1}^{n}$. For simplicity, we have ignored the enforcement of cluster sizes (otherwise we need to find a pair of such nodes). To proceed, we estimate the probability mass function (PMF) of $\tilde{N}^{(i)}_{\text{inter}}$ as, for $k\geq 1$,
$\mathcal{P}(\tilde{N}^{(i)}_{\text{inter}}=k)$ $=\mathcal{P}\left(g_{ij}=g_{i}g^{-1}_{j},\;\forall j\text{ connected to }i,\text{ and }\kappa_{j}\neq\kappa_{i}\mid N^{(i)}_{\text{inter}}=k\right)\cdot\mathcal{P}\left(N^{(i)}_{\text{inter}}=k\right)$
$\overset{(a)}{\approx}\frac{1}{M^{k}}\cdot{n/2\choose k}q^{k}(1-q)^{\frac{n}{2}-k}\approx\frac{e^{-\frac{nq}{2}}}{k!}\left(\frac{nq}{2M}\right)^{k}.$ (15)
In the last step of (15), we apply the Poisson approximation that $\text{Binom}\left(\frac{n}{2},q\right)\approx\text{Poisson}\left(\frac{nq}{2}\right)$. Notably, in step $(a)$ we approximate that the probability for $k$ inter-cluster edges to satisfy the group transformation consistency is $1/M^{k}$, which aligns with our model setting where each edge is drawn uniformly at random from $M$ group elements. However, we emphasize that this assumption requires careful justification that is provided in Section 6.3, which plays a critical role in our proof. Given the PMF (15), the moment generating function (MGF) of $\tilde{N}^{(i)}_{\text{inter}}$ can be approximated as
$\mathbb{E}\left[e^{t\tilde{N}^{(i)}_{\text{inter}}}\right]$ $=\underbrace{1-\sum_{k=1}^{\infty}\mathcal{P}(\tilde{N}^{(i)}_{\text{inter}}=k)}_{\mathcal{P}(\tilde{N}^{(i)}_{\text{inter}}=0)}+\sum_{k=1}^{\infty}\mathcal{P}(\tilde{N}^{(i)}_{\text{inter}}=k)e^{tk}\approx 1-\sum_{k=1}^{\infty}\frac{e^{-\frac{nq}{2}}}{k!}\left(\frac{nq}{2M}\right)^{k}\cdot(1-e^{tk})$
$=1-\exp\left(-\left(1-\frac{1}{M}\right)\frac{nq}{2}\right)+\exp\left(-\left(1-\frac{e^{t}}{M}\right)\frac{nq}{2}\right).$
Applying the Chernoff bound, we have, for any $t>0$
$\mathcal{P}\left(\tilde{N}^{(i)}_{\text{inter}}\geq N^{(i)}_{\text{inner}}\right)\leq\mathbb{E}\left[e^{t(\tilde{N}^{(i)}_{\text{inter}}-N^{(i)}_{\text{inner}})}\right].$ (16)
As $\tilde{N}^{(i)}_{\text{inter}}$ and $N^{(i)}_{\text{inner}}$ are independent, we can further bound the RHS of (16) as
$\log\left(\mathbb{E}\left[e^{t(\tilde{N}^{(i)}_{\text{inter}}-N^{(i)}_{\text{inner}})}\right]\right)$ ${=}\log\left(\mathbb{E}\left[e^{t\tilde{N}^{(i)}_{\text{inter}}}\right]\right)+\log\left(\mathbb{E}\left[e^{-tN^{(i)}_{\text{inner}}}\right]\right)$
$\overset{(a)}{\approx}\log\left(1-\exp\left(-\left(1-\frac{1}{M}\right)\frac{nq}{2}\right)+\exp\left(-\left(1-\frac{e^{t}}{M}\right)\frac{nq}{2}\right)\right)+\frac{np}{2}\left(e^{-t}-1\right)$
$\overset{(b)}{\approx}\frac{n}{2}\left(q\left(\frac{e^{t}}{M}-1\right)+p\left(e^{-t}-1\right)\right),$ (17)
where $(a)$ applies the Poisson approximation on $N^{(i)}_{\text{inner}}$ similar to $\tilde{N}^{(i)}_{\text{inter}}$ in (15), and $(b)$ holds due to $e^{-t}<M$ for $t>0$, which will indeed be the case when optimizing with regard to $t$ in (18), followed by using the approximations $e^{x}\approx 1+x$ and $\log(1+x)\approx x$. Now, as $t$ can be chosen arbitrary, by taking
$t=\frac{1}{2}\log\left(\frac{Mp}{q}\right)=\frac{1}{2}\log\left(\frac{Ma}{b}\right),$ (18)
to minimize the RHS of (17), we obtain
$\log\left(\mathbb{E}\left[e^{t(\tilde{N}^{(i)}_{\text{inter}}-N^{(i)}_{\text{inner}})}\right]\right)\leq-\left(\frac{a+b}{2}-\sqrt{\frac{ab}{M}}\right)\log n.$
By plugging this into (16), we have
$\mathcal{P}\left(\tilde{N}^{(i)}_{\text{inter}}>N^{(i)}_{\text{inner}}\right)\leq\exp\left(-\left(\frac{a+b}{2}-\sqrt{\frac{ab}{M}}\right)\log n\right).$
Applying the union bound over all the $n$ nodes yields
[…]
Therefore, the failure probability goes to 00 when $\frac{a+b}{2}-\sqrt{\frac{ab}{M}}>1$.
We remark that the above analysis is based on several approximations and assumptions. For instance, we do not enforce the cluster size constraint, and the consistency analysis is entirely omitted in (15). Nevertheless, this simplified argument yields the same parameter threshold as the more rigorous analysis presented in Section 6. As such, it serves as a warm-up proof and provides intuitive justification for the theorem.
@@ 3.2 Exact recovery of group elements
then the MLE (1) exactly recovers the group elements w.p. at least $1-n^{-c+o(1)}$, i.e.,$\mathcal{P}_{e,\mathrm{g}}(\psi_{\textup{MLE}})\leq n^{-c+o(1)}$.
@@
then the MLE (1) fails to exactly recover the group elements w.p. at least $0.45$, i.e.,$\mathcal{P}_{e,\mathrm{g}}(\psi_{\textup{MLE}})\geq 0.45$.
The proofs of Theorems 3.3 and 3.4 are deferred to Section 8 and Section 9 respectively, which are simply extended from the ones of Theorems 3.1 and 3.2. Similar to the threshold (10), Theorems 3.3 and 3.4 suggest a phase transition exists for the exact recovery of group elements, given as